OpenStudy (anonymous):

What is the decimal equivalent to \frac{1}{20}? A. 0.5 B. 0.2 C. 0.05 D. 0.02

OpenStudy (anonymous):

this is what calculators are made for

OpenStudy (alexandervonhumboldt2):

@satellite73 this is a mind problem

OpenStudy (anonymous):

@satellite73 how rude!

OpenStudy (anonymous):

although you could a) divide or b) write \(f\rac{1}{20}\times \frac{5}{5}=\frac{5}{100}\)

OpenStudy (fibonaccichick666):

ok, so for this style, you need to either 1. Do the long division. Or two, change your fraction to something over 100

OpenStudy (anonymous):

oops

OpenStudy (alexandervonhumboldt2):

1/20 is too easy to use a calculator: 100/20=5. multply the 1/20 * 5/5=5/100=0.05

OpenStudy (anonymous):

Calculators? That's nice.. Don't be so rude. I don't know the slightest things about fractions. And you expect me to do something with a calculator. Thanks.

OpenStudy (anonymous):

\[\frac{1}{20}\times \frac{5}{5}=\frac{5}{100}\]

OpenStudy (anonymous):

i was totally serious calculators are made for decimals' put in \(1\div 20\) and you will get the decimal answer

OpenStudy (fibonaccichick666):

A fraction is just a division problem. If you can do long division, that is one way.

OpenStudy (alexandervonhumboldt2):

but you will not always have a calculator by your hand...

OpenStudy (anonymous):

I didn't know that xD

OpenStudy (anonymous):

check it really the fraction bar means "divided by" so for example \[\frac{1}{8}=1\div 8=.125\]

OpenStudy (fibonaccichick666):

the other option is to make the denominator a multiple of ten. When you do this, you can directly relate the value to the place after the decimal. Just like we did with 625/10000

OpenStudy (anonymous):

I got 0.05 in the claculator. Confirmation?

OpenStudy (anonymous):

yes

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (fibonaccichick666):

that is true, though you could still do it by hand

OpenStudy (anonymous):

don't be confused by these fractions it means division, that is all most fractions make lousy decimals (try \(\frac{1}{7}\) and you will see why)

OpenStudy (anonymous):

Fractions are a pain in the bleep. Anymore with my scvhool ;-;

OpenStudy (fibonaccichick666):

well, if you ant another way to look at it, they are just multiplication of an inverse XD

OpenStudy (anonymous):

yeah they are a pain to work with, but pretty necessary for computing and measuring stuff

OpenStudy (fibonaccichick666):

no such thing as division in my book

OpenStudy (anonymous):

lol

OpenStudy (fibonaccichick666):

no such thing as subtraction either. There is only addition. then, we get sick of writing addition so we have multiplication